OpenStudy (moonlitfate):

Consider the parametric equations below. x = 1 + 2t, y = 4 − t^2 Eliminate the parameter to find a Cartesian equation of the curve.

OpenStudy (dumbcow):

solve one equation for "t" , then substitute it into other equation.

OpenStudy (dumbcow):

\[x = 1+2t \rightarrow t = \frac{x-1}{2}\] \[\rightarrow y = 4 - (\frac{x-1}{2})^2\]